a series rlc circuit as a resistance of 6.0 ohm capacitenace of 5882 mf and inductance of 100mH. what is the y location in rad/sec of the pole s1?
1
Expert's answer
2020-10-19T08:15:11-0400

The "y" pole position "S_1" will be equal to the cyclic frequency "\omega =\frac {1} {\sqrt {L\times C}} =\frac {1} {\sqrt {100\times 10 ^ {-3 }\times 5882\times 10 ^ {-3}}}=1.3"  rad/sec.